OpenStudy (anonymous):

Please help. Easy medal. Find an equation for the line passing through the origin and parallel to the line passing through the points (2,4) and (4,7) Please.

OpenStudy (anonymous):

Wait a sec

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Okay then let me get this straight there are two lines one of which passes through the origin and the other passes through the coordinates mentioned?

OpenStudy (anonymous):

Yes, correct.

OpenStudy (anonymous):

First you find the gradient of the line passing through the coordinates mentioned

OpenStudy (anonymous):

This gradient is going to be the same for the line passing through the origin

OpenStudy (anonymous):

I have never learned about a gradient. Please guide me through that?

OpenStudy (anonymous):

Do you know what a slope is?

OpenStudy (anonymous):

Do you mean slope?

OpenStudy (anonymous):

Slope=gradient

OpenStudy (anonymous):

Of course. Yes.

OpenStudy (anonymous):

Let me do that first.

OpenStudy (anonymous):

Then you find the slope of the line passing through the coordinates mentioned

OpenStudy (anonymous):

You will get 3/2

OpenStudy (anonymous):

yes. I got 3/2

OpenStudy (anonymous):

Nice then the next step This gradient is equal for the line passing through the origin as these lines are parallel

OpenStudy (anonymous):

0=3/2x0+c

OpenStudy (anonymous):

You dont need to do the above step but i still did

OpenStudy (anonymous):

Oh, I see. Then, I believe I use the slope-intercept formula for the next coordinates? Like y-y1= m(x-x1)? Or..is that incorrect?

OpenStudy (anonymous):

So the equation of line should be y=3x/2+0

OpenStudy (anonymous):

Just substitute the m and the (0,0) to get the c which we already know is 0 as the line passes through the origin you got your equ.

OpenStudy (anonymous):

Ohhh...okay, I understand now.

OpenStudy (anonymous):

You can use that formula

OpenStudy (anonymous):

I guess it can work, but the way you explained it since the origin is only (0,0) and the slope off of the second coordinates is 3/2 makes it pretty obvious. Thanks very much for your help

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

Good luck

OpenStudy (anonymous):

Thanks.